OpenStudy (anonymous):

f(x) = (2x-3)/(x+4). find f^-1(x)

Parth (parthkohli):

\[f^{-1}(x)\]?

OpenStudy (anonymous):

yes

hartnn (hartnn):

where r u stuck?

hartnn (hartnn):

i know that u know the procedure

Parth (parthkohli):

OK,\[y = f(x)\]\[y = {2x - 3 \over x +4}\]Plug in \(x\) with \(y\).\[x = {2y - 3 \over y + 4}\]Solve for \(x\).

OpenStudy (anonymous):

well i worked it out and got a quadratic on one side.. so now i started over.

OpenStudy (anonymous):

@ParthKohli if you plug in x with y .. would you then solve for y instead of x?

OpenStudy (anonymous):

so after multiplying y+4 to both sides.. i get xy+4x = 2y-3 am i on the right track?

hartnn (hartnn):

\(\huge x=\frac{2y+8-11}{y+4}=\frac{2(y+4)}{y+4}-\frac{11}{y+4}\)

hartnn (hartnn):

=2-11/(y+4) subtract 2 from both sides.

hartnn (hartnn):

did u get that step ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

my paper is a mess now :( i got the inverse as x+5.. sigh

OpenStudy (anonymous):

\[x=\frac{ 2y-3 }{y+4 }\]

OpenStudy (anonymous):

then i multiply both sides by Y+4 to get

hartnn (hartnn):

don't do that

OpenStudy (anonymous):

xy+4x = 2y -3

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

don't we just take the derivative and plug in 1?

OpenStudy (anonymous):

i dont even know what that means :)

hartnn (hartnn):

write -3 as 8-11 2y-3 = 2y+8-11 = 2(y+4) -11

hartnn (hartnn):

i don't think u are getting this ^ @Nameless

OpenStudy (anonymous):

i am trying to derive a way to do every problem like this.... thus if it is x = y-1/3x+5

OpenStudy (anonymous):

then i simply replace x with y and solve for y. multiply both sides by 3y+5 then rearrange and factor out the y to get y(3x-1) = -5x-1 .. then divide 3x-1 by -5x-1 to get y = -5x-1/3x-1

OpenStudy (anonymous):

i am trying to do it your way.. but its different than i have learned in class.. so

hartnn (hartnn):

ok, u have x= (2y-3) / (y+4) x(y+4) = 2y-3 xy+4x = 2y-3 xy-2y+4x = -3 xy -2y =-3-4x y(x-2) = -(3+4x) y=-(3+4x)/ (x-2) make sense ?

OpenStudy (anonymous):

yes!

hartnn (hartnn):

ask if any doubts in any step ...

OpenStudy (anonymous):

the one u typed is the one i had been doing.. but got confused with the algebra... mostly what to move.. i see now, thank you so very much

hartnn (hartnn):

welcome :)